# Enhanced Histopathology Image Feature Extraction using EfficientNet with Dual Attention Mechanisms and CLAHE Preprocessing

### v3

@@ Abstract
Breast cancer histopathology image classification is critical for early detection and improved patient outcomes. This study introduces a novel approach leveraging Hybrid EfficientNet models enhanced with advanced attention mechanisms—CBAM, Self-Attention, and Deformable Attention—to improve feature extraction and focus on relevant tissue regions. The proposed models were evaluated on the BreakHis dataset across multiple magnification scales (40X, 100X, 200X, and 400X). Among them, the EfficientNetV2-XL with CBAM achieved outstanding performance, reaching a peak accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, outperforming state-of-the-art methods. By integrating Contrast Limited Adaptive Histogram Equalization (CLAHE) for preprocessing and optimizing computational efficiency, this method demonstrates its suitability for real-time clinical deployment. The results underscore the potential of attention-enhanced scalable architectures in advancing diagnostic precision for breast cancer detection.
@@
Transfer Learning, Modality Specific, EfficientNet, CBAM, Self-Attention, Deformable Attention, Histopathology, Image Classification, Deep Learning, Breast Cancer.
@@ I Introduction
Breast cancer remains one of the most prevalent and deadly forms of cancer, especially affecting women worldwide. According to global cancer statistics, breast cancer accounts for more than 2.3 million new cases each year, contributing to a significant number of cancer-related deaths [3]. Early detection plays a critical role in reducing the mortality rate associated with this disease. The standard diagnostic procedure for detecting breast cancer involves histopathological analysis, where biopsy tissue samples are examined under a microscope. This process is time-consuming, labor-intensive, and prone to subjective interpretation errors, even among trained pathologists [1].
Pathologists rely on morphological abnormalities in cell structures, particularly the nuclei, to distinguish between benign and malignant cells [2]. Benign tumors typically exhibit well-organized cellular structures and rarely invade surrounding tissues or metastasize to other parts of the body. In contrast, malignant tumors display uncontrolled growth, irregular nuclei, and the potential to spread, making them life-threatening. Early identification of these malignant traits is critical for improving patient prognosis and guiding treatment plans.
In recent years, computer-aided diagnosis (CAD) systems have been developed to assist pathologists by automating the analysis of histopathology images. Deep learning (DL) models, particularly Convolutional Neural Networks (CNNs), have shown great promise in this domain due to their ability to automatically learn complex features from medical images [7]. However, despite their success in natural image classification, traditional CNN architectures, such as VGGNet and ResNet, struggle to capture the intricate features present in histopathological images, primarily due to the variability in magnification scales and the complex structure of the tissue samples.
To address these challenges, researchers have begun exploring hybrid models that combine CNNs with attention mechanisms. These attention mechanisms, including the Convolutional Block Attention Module (CBAM), Self-Attention, and Deformable Attention, have been designed to enhance feature extraction by allowing the network to focus on the most relevant parts of the image. CBAM, for example, has shown considerable potential in medical image analysis by directing the network’s attention toward critical regions, leading to improved classification accuracy. Similarly, Self-Attention mechanisms, initially introduced in transformer models for Natural Language Processing (NLP), have been adapted to medical imaging tasks, where they capture long-range dependencies across the image. Deformable Attention further enhances this by dynamically adjusting the receptive field based on the complexity of the input features, making it particularly effective for identifying abnormalities in histopathology images.
In this study, we propose a series of Hybrid EfficientNet models integrated with advanced attention mechanisms to improve the performance of breast cancer histopathology image classification. EfficientNet, a highly scalable CNN architecture, has been recognized for its ability to achieve state-of-the-art performance with fewer parameters by optimizing the depth, width, and resolution of the model. By incorporating attention mechanisms such as CBAM, Self-Attention, and Deformable Attention into EfficientNet, we aim to enhance the network’s focus on key features, thereby improving the classification of benign and malignant breast cancer tissue samples.
The dataset used in this study is the BreakHis dataset, a publicly available dataset widely used for breast cancer histopathology image classification tasks . The dataset consists of 9,109 images at four different magnification levels (40X, 100X, 200X, and 400X), providing a comprehensive set of data to test the proposed models across varying scales of tissue detail. The performance of the proposed models is evaluated using key metrics such as accuracy, precision, recall, and F1-score.
The rest of the paper is structured as follows: In Section II, we review the existing literature on breast cancer histopathology image classification and attention mechanisms. Section III describes the methodology, including the dataset, preprocessing techniques, model architecture, and training procedures. Section IV presents the experimental results and provides a discussion on the comparative performance of the models. Finally, Section V concludes the paper and outlines potential directions for future research.
Automated histopathology image classification has seen substantial improvements with the advent of deep learning techniques, particularly with convolutional neural networks (CNNs). However, existing methods, such as using generic CNNs like VGGNet and ResNet, may fail to capture critical features from histopathology images due to their complex structure and variability in magnification scales. EfficientNet models have demonstrated strong performance in image classification tasks due to their scalability and efficiency. To further enhance feature extraction, attention mechanisms such as Convolutional Block Attention Module (CBAM), Self-Attention, and Deformable Attention have been proposed.
In this study, we explore Hybrid EfficientNet models that incorporate these advanced attention mechanisms to address the challenges posed by histopathological image variability. The models are evaluated across multiple magnification scales (40X, 100X, 200X, 400X) using the BreakHis dataset, which is a widely recognized dataset for breast cancer histopathology image classification.
@@ II Literature Review
In recent years, extensive research has been conducted on breast cancer classification using deep learning, particularly convolutional neural networks (CNNs). Several studies have explored the use of hybrid models and attention mechanisms to improve the performance of CNNs in histopathology image classification.
Chuang Zhu et al. proposed a hybrid model that assembles multiple compact CNNs for breast cancer classification. This architecture includes a squeeze excitation pruning block and local and global branches, which help reduce channel redundancy and provide a strong representation of the input images. Experiments were performed using the BreakHis and BACH datasets, and the multi-model assembling strategy achieved comparable results with state-of-the-art models Another approach by Togacar et al. developed a residual architecture called BreastNet, which includes attention modules for identifying key regions in histopathology images. The hypercolumn technique was also employed to enhance feature representation. The BreastNet model achieved an accuracy of 98.80% using dense pooling and residual blocks.
Similarly, EM Nejad et al. introduced a single-layer CNN to extract salient features from histopathology images from the BreakHis dataset. Their model achieved a recognition rate of 77.5% . AA Nahid et al. combined Long Short-Term Memory (LSTM) networks with CNNs for breast cancer classification, using Support Vector Machines (SVM) and softmax for final decision-making. Their model achieved 91.00% accuracy.
Yan et al. proposed a model that integrates recurrent neural networks (RNN) with CNNs for histopathology image analysis. Using the BreakHis dataset, they achieved an accuracy of 91.3% for the classification of four cancer classes. Xie et al. also explored deep convolutional neural networks (DCNN) for binary and multi-class breast cancer classification. Their approach used InceptionResNetV2 to extract features, which were further compressed using an autoencoder. This model achieved improved clustering results.
Several studies have leveraged transfer learning to mitigate the data scarcity issue in medical imaging. Sana Ullah et al. developed a model that extracts features from pre-trained networks like VGGNet, GoogleNet, and ResNet. These features are then passed through fully connected layers for binary classification, achieving 97.525% accuracy. Ahmad et al. applied transfer learning using AlexNet, ResNet, and GoogleNet for breast cancer classification, obtaining a maximum accuracy of 85% with ResNet . In another study, Vesal et al. applied transfer learning to classify breast cancer into four sub-classes using the BACH dataset. They utilized ResNet50 and InceptionV3, achieving an accuracy of 97.50% .
Despite these advances, many of the existing methods fail to fully capture the complex structure of breast histopathology images, particularly at different magnification scales. This has led to the exploration of attention mechanisms, which improve the ability of models to focus on the most relevant regions of the image. Woo et al. proposed the Convolutional Block Attention Module (CBAM), which applies attention along both the channel and spatial dimensions to enhance feature extraction . By focusing on critical areas, CBAM has demonstrated significant performance improvements in medical image classification tasks.
In this context, our work aims to further enhance breast cancer classification by integrating attention mechanisms, such as CBAM, Self-Attention, and Deformable Attention, with the EfficientNet architecture. EfficientNet has demonstrated state-of-the-art performance in image classification tasks due to its compound scaling strategy, which optimally balances model depth, width, and resolution.
Our proposed approach builds upon existing research by incorporating advanced attention mechanisms and leveraging the strengths of EfficientNet, making it more robust in handling the variability of histopathology images. Through comprehensive experimentation and comparison with other state-of-the-art methods, we aim to demonstrate that our Hybrid EfficientNet models significantly improve breast cancer classification accuracy, particularly at higher magnification scales.
@@ III Methodology
III-A Hybrid EfficientNet Model Architecture
The backbone of our proposed model consists of the EfficientNet architecture, specifically EfficientNetV2-XL and EfficientNetV2-L. These models leverage compound scaling to optimize accuracy and efficiency, achieving state-of-the-art results in various image classification tasks. The architecture utilizes a combination of depth, width, and resolution scaling to adapt to the complexities of histopathological images, which often exhibit considerable variability in cellular structures and staining quality.
To enhance feature extraction capabilities, we integrate the following advanced attention mechanisms:
  • Convolutional Block Attention Module (CBAM): The CBAM module consists of two sequential attention mechanisms—channel attention and spatial attention. The channel attention mechanism allows the model to focus on informative features by leveraging global information from feature maps, while spatial attention helps the model concentrate on salient regions within the feature maps. This two-step process is crucial for enhancing the network’s focus on pertinent structures in histopathological images.
  • Self-Attention: The Self-Attention mechanism captures long-range dependencies across the input image, enabling the model to aggregate contextual information from various parts of the image. This is particularly important for histopathology images, where cancerous patterns may be dispersed throughout the tissue.
  • Deformable Attention: This mechanism adaptively adjusts the receptive fields in response to the input features, allowing the model to better capture complex shapes and variations common in histopathological images. The deformable attention layers are instrumental in adapting to the irregularities of tumor shapes and sizes, thus improving classification accuracy.
The integrated architecture, depicted in Fig. 1, begins with the EfficientNet backbone, followed by attention modules that refine the feature representation before passing through fully connected classification layers.
III-B CBAM-EfficientNet Architecture
@@ III-B3 EfficientNet Backbone
This combined approach of CBAM and EfficientNet leverages both channel and spatial attention mechanisms, providing a robust framework for the classification of breast histopathology images while maintaining computational efficiency.
@@ III-E Training and Hyperparameter Tuning
The training of the models was conducted using PyTorch on an NVIDIA RTX 4060 GPU, enabling efficient computation. We utilized an initial learning rate of 0.001 and employed both the Adam and SGD optimizers to determine which provided better convergence and generalization.
@@
  • Activation Functions: Both ReLU and tanh activation functions were tested at the last classification layer to evaluate their impact on the model’s performance.
@@
  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks.
  • Batch Size and Epochs: We tested various batch sizes (16, 32, and 64) and trained the models for a maximum of 100 epochs, applying early stopping based on validation loss to prevent overfitting.
III-F Modality-Specific Transfer Learning Strategy
@@ III-G Evaluation Metrics
Figure 4: ROC-AUC Curve of EfficientNetV2-XL with CBAM
TABLE II: Performance of Various Models Across Magnification Scales
Model Scale Accuracy (%) F1 Score (%)
EfficientNetV2-XL + CBAM 40X 95.02 94.13
100X 96.85 95.69
200X 97.45 96.96
400X 98.96 98.31
EfficientNetV2-L + CBAM 40X 94.89 97.78
100X 96.13 97.99
200X 97.85 98.03
400X 99.16 98.15
EfficientNetV2-XL + Self-Attention 40X 96.56 92.02
100X 97.35 93.29
200X 97.13 94.16
400X 98.56 95.43
EfficientNetV2-L + Self-Attention 40X 96.23 91.89
100X 97.13 92.78
200X 97.89 93.93
400X 98.45 94.89
EfficientNetV2-XL + Deformable Attention 40X 96.35 95.89
100X 96.89 96.35
200X 97.13 96.56
400X 97.56 96.93
EfficientNetV2-L + Deformable Attention 40X 96.13 94.17
100X 96.78 94.23
200X 96.89 95.35
400X 97.45 96.49
IV Experimental Results and Discussions
IV-A Experimental Platform
The experiments were conducted using PyTorch as the deep learning framework backend. The models were trained and tested on an NVIDIA RTX 4060 GPU, providing substantial computational power for the deep learning tasks involved. The system was equipped with 16GB of RAM to efficiently handle the large dataset and facilitate faster processing.
IV-B Hyperparameter Optimization
@@
We evaluated two different optimizers: Adam (Adaptive Moment Estimation) and SGD (Stochastic Gradient Descent), with three distinct learning rates ($0.01$, $0.001$, and $0.0001$). The SoftMax classifier was employed for all classification experiments, providing probability scores for each class label.
@@
IV-C Results
We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table III.
@@ IV-C Results
In contrast, the EfficientNetV2-XL + CBAM was trained on a more modest NVIDIA RTX 4060 GPU with just 8GB of VRAM. Despite the lower memory availability, the EfficientNetV2-XL architecture demonstrated exceptional computational efficiency, leveraging depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM). This efficiency allowed the model to maintain superior accuracy and scalability, particularly at higher magnifications, while requiring significantly fewer computational resources.
TABLE III: Comparison of Accuracy Across Models and Magnification Factors
@@ V Conclusion
The Hybrid EfficientNetV2-XL + CBAM and previous SOTA VGG16 + VGG19 model achieved superior accuracy, however EfficientNetv2-XL has F1-Score of 98.31 and Hybrid VGG16 + VG199 has 97.95% ,therefore EfficientNetV2-XL + CBAM outperforms other models in both accuracy and computational efficiency. By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics. These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models, highlighting its potential for real-time clinical deployment. Future work may adapt the model to other datasets and further expand its clinical applicability.

### v4

# CBAM-EfficientNetV2 for Histopathology Image Classification using Transfer Learning and Dual Attention Mechanisms

@@ Abstract
Breast cancer histopathology image classification is critical for early detection and improved patient outcomes. This study introduces a novel approach leveraging EfficientNetV2 models, to improve feature extraction and focus on relevant tissue regions. The proposed models were evaluated on the BreakHis dataset across multiple magnification scales (40X, 100X, 200X, and 400X). Among them, the EfficientNetV2-XL with CBAM achieved outstanding performance, reaching a peak accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, outperforming state-of-the-art methods. By integrating Contrast Limited Adaptive Histogram Equalization (CLAHE) for preprocessing and optimizing computational efficiency, this method demonstrates its suitability for real-time clinical deployment. The results underscore the potential of attention-enhanced scalable architectures in advancing diagnostic precision for breast cancer detection.
@@
Transfer Learning, Modality Specific, EfficientNet, CBAM, Histopathology, Image Classification, Deep Learning, Breast Cancer.
@@ I Introduction
Automating histopathological analysis has become a vital research area in medical imaging due to its potential to significantly reduce the diagnostic workload of pathologists and enhance the consistency and reproducibility of results. Histopathology—the microscopic study of tissue abnormalities—is fundamental in diagnosing diseases such as cancer, but manual analysis is labor-intensive, subjective, and prone to variability between observers. As a result, computer-aided diagnosis (CAD) systems have gained traction, with the aim of augmenting pathologists’ abilities to efficiently and accurately classify tissue samples [1]. Despite substantial progress, CAD systems continue to face significant challenges, particularly in handling the variability inherent in histopathological images. Differences in magnification levels, staining protocols (e.g., Hematoxylin and Eosin, or H&E staining), and tissue morphology introduce variability that complicates the generalization of standard deep learning models [2].
Traditional convolutional neural networks (CNNs), while successful in many computer vision tasks, often struggle to generalize across these variations in histopathological data. Specifically, CNNs are designed with fixed receptive fields and are limited in their ability to capture fine-grained details that vary across multiple scales [3]. The heterogeneity of breast cancer histopathology images—marked by different magnification levels and complex cellular structures—poses unique challenges that demand more sophisticated architectures. To address this, researchers have turned to hybrid models that integrate attention mechanisms with CNNs, which have demonstrated superior performance in medical imaging tasks by enhancing feature extraction.
Attention mechanisms such as the Convolutional Block Attention Module (CBAM) have been adapted to improve image classification by enabling the network to focus dynamically on the most informative regions of an image. In histopathology, these mechanisms are particularly effective in highlighting regions critical to cancer diagnosis, such as irregular nuclei shapes, abnormal cellular arrangements, and other key tissue structures that may indicate malignancy. By applying attention, the network can dynamically recalibrate feature maps along spatial and channel dimensions, improving its ability to capture both local and global image features [4]. This is especially valuable in distinguishing between benign and malignant tissue, where small but significant differences in morphology must be detected.
EfficientNet [8], a state-of-the-art CNN architecture known for its compound scaling strategy, is particularly well-suited for medical image classification tasks due to its ability to balance model depth, width, and resolution. EfficientNet achieves superior performance with fewer parameters, making it computationally efficient without sacrificing accuracy. By integrating attention mechanisms with EfficientNet, we aim to build models that can more effectively classify breast cancer histopathology images while remaining scalable to real-world clinical applications. This hybrid approach addresses the need for models that can operate across varying image resolutions and magnification levels, ensuring robustness and high accuracy in diverse settings.
@@ II Literature Review
Hybrid models have gained significant attention in breast cancer classification due to their ability to combine the strengths of multiple architectures. A multi-model assembling strategy that integrates local and global branches, reducing redundancy and enhancing feature representation. This approach has achieved high accuracy on standard breast cancer datasets, demonstrating the potential of hybrid architectures. Similarly, Togacar et al. [5] developed BreastNet, a model that employs attention mechanisms along with hypercolumn techniques to improve the precision of identifying malignant regions. The inclusion of attention modules has been shown to substantially improve performance, especially in complex tasks like cancer detection.
The use of attention mechanisms in medical imaging has proven highly effective. Woo et al. [4] introduced CBAM, which applies sequential attention along both the channel and spatial dimensions. This method allows networks to highlight critical regions in complex medical images, contributing to substantial performance improvements in classification tasks involving histopathology images. Beyond CBAM, Self-Attention [6] has been adapted from its initial use in natural language processing to capture contextual dependencies across different regions of an image. This has proven particularly useful in medical image analysis, where understanding the relationship between local and global image features is essential for accurate classification.
Recent advancements in hybrid attention mechanisms have significantly improved the capability of deep learning architectures to focus on salient features in complex datasets. For instance, works like Transformer-CNN hybrids [6] have shown remarkable success in combining global context capture through self-attention with local feature extraction inherent in CNNs, making them highly effective for tasks such as lesion segmentation and tissue classification in histopathology.
Another notable development is the Multi-Scale Attention Networks (MSANs), which leverage hierarchical attention mechanisms to capture fine-grained details and high-level contextual information simultaneously. These architectures have proven particularly effective for analyzing multi-resolution data, as demonstrated by Zhang et al. [35], who employed MSANs for high-resolution cellular analysis in breast cancer histopathology, achieving state-of-the-art performance.
Furthermore, Deformable Attention Mechanisms have recently gained traction for their adaptability in handling heterogeneous medical datasets. He et al. [36] introduced deformable attention layers that dynamically adjust receptive fields, outperforming traditional attention mechanisms in scenarios with high inter-patient variability, such as BreakHis and CAMELYON17 datasets.
Hybrid approaches like these emphasize the potential of combining localized feature emphasis (via channel/spatial attention) with global dependency modeling, bridging a critical gap in medical imaging where both local abnormalities and global patterns must be identified. Our approach, integrating CBAM with EfficientNetV2, aligns with this trend but distinguishes itself by optimizing computational efficiency and scalability for clinical deployment, as evidenced by its performance on the BreakHis dataset.
@@ III Methodology
III-A Model Architecture
Our proposed model is based on EfficientNetV2-XL which use compound scaling to optimize accuracy and efficiency, making them ideal for handling the variability in histopathological images, such as differences in cellular structures and staining quality.
To enhance feature extraction, we integrate the Convolutional Block Attention Module (CBAM). CBAM combines channel and spatial attention, allowing the model to focus on informative features and key regions within the image. This is particularly beneficial in histopathology, where precise identification of cancerous structures is crucial. The EfficientNet backbone extracts initial features, and CBAM refines these representations before passing through fully connected layers for classification.
This integrated architecture, illustrated in Fig. 1, optimizes both efficiency and accuracy in breast cancer histopathology classification.
III-B CBAM Architecture
@@ III-B3 EfficientNet Backbone
This combined approach of CBAM and EfficientNet leverages both channel and spatial attention mechanisms, providing a robust framework for the classification of breast histopathology images while maintaining computational efficiency
@@ III-E Training and Hyperparameter Tuning
The training of the models was conducted using PyTorch on CUDA, enabling efficient computation. We utilized an initial learning rate of 0.001 and employed Adam optimizers to determine which provided better convergence and generalization.
@@
  • Activation Functions:ReLU activation functions was used at the last classification layer.
@@
  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks with a rate of 0.4.
  • Batch Size and Epochs: We tested various batch sizes (16, 32, and 64) and finalized with 32 and trained the models for a maximum of 100 epochs, applying early stopping based on validation loss to prevent overfitting.
@@ III-F Evaluation Metrics
The experiments were conducted using PyTorch as the deep learning framework backend. The models were trained and tested on an NVIDIA RTX 4080 GPU, providing substantial computational power for the deep learning tasks involved. The system was equipped with 16GB of RAM to efficiently handle the large dataset and facilitate faster processing.
@@
We evaluated two different optimizers: Adam (Adaptive Moment Estimation), with learning rate ( $0.001$. The SoftMax classifier was employed for all classification experiments, providing probability scores for each class label.
@@
Figure 6: ROC-AUC Curve of EfficientNetV2-XL with CBAM
IV Experimental Results and Discussions
IV-A Results
We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table II.
@@ IV-A Results
In contrast, the EfficientNetV2-XL + CBAM was trained on a more modest NVIDIA RTX 4080 GPU with just 16GB of VRAM. Despite the lower memory availability, the EfficientNetV2-XL architecture demonstrated exceptional computational efficiency, leveraging depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM).
TABLE II: Comparison of Accuracy Across Models and Magnification Factors
@@ V Conclusion
The Hybrid EfficientNetV2-XL + CBAM and previous State-of-the-Art model VGG16 + VGG19 model achieved superior accuracy, however EfficientNetv2-XL has F1-Score of 98.31 and Hybrid VGG16 + VG199 has 97.95% ,therefore EfficientNetV2-XL + CBAM outperforms other models in both accuracy and computational efficiency. By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics. These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models, highlighting its potential for real-time clinical deployment. Future work may adapt the model to other datasets and further expand its clinical applicability.